OpenStudy (anonymous):

limits! limit of sqrt(x+3)-sqrt(3x+1)/x-1 as x approaches 1

OpenStudy (anonymous):

\[\lim_{x \rightarrow1} \frac{ \sqrt{x+3}-\sqrt{3x+1} }{x-1 }\]

OpenStudy (anonymous):

rationalize

OpenStudy (anonymous):

\[ L=\lim_{x\to1}\frac{ \sqrt{x+3}-\sqrt{3x+1} }{x-1}\times\frac{\sqrt{x+3}+\sqrt{3x+1}}{\sqrt{x+3}+\sqrt{3x+1}} \]

OpenStudy (anonymous):

\[L=\lim_{x\to1}\frac{(x+3)-(3x+1)}{(x-1)(\sqrt{x+3}+\sqrt{3x+1})}\]

OpenStudy (anonymous):

ok and then do you plug in 1 for x?

OpenStudy (anonymous):

can you? ;)

OpenStudy (anonymous):

uh no...that gives me 0 on the denominator

OpenStudy (anonymous):

exactly.. what was the point of doing that extra multiplication? .. simplify that numerator a little bit more

OpenStudy (anonymous):

\[x+3-3x-1=-2x+2=-2(x-1)\]

OpenStudy (anonymous):

got it! -1/2 thank you:)

OpenStudy (anonymous):

yep.